OpenStudy (anonymous):

Medal Find the GCF. 40, 72

OpenStudy (princeharryyy):

fig???

OpenStudy (anonymous):

what?

OpenStudy (princeharryyy):

figure

OpenStudy (anonymous):

there is no figure

OpenStudy (princeharryyy):

sorry the answer isn't possible.. then, still try 180 - 40 - 72 = 68 if it is a triangle

OpenStudy (princeharryyy):

@harz360

OpenStudy (usukidoll):

lol use gcd( 40,72) and long division multiple times until you reach gcd(0,1) and write it as a linear combination of lalalallallalalalllallalal

OpenStudy (anonymous):

i dont understand

OpenStudy (princeharryyy):

oh my god i guess u r right @UsukiDoll that meant gcf.. n tnks

OpenStudy (usukidoll):

I'm using number theory on this XD

OpenStudy (princeharryyy):

greatest common factor

OpenStudy (anonymous):

okay

OpenStudy (usukidoll):

gcd and gcf mean the same thing... if you know number theory and just divide so many times and put them back as a linear combination wham. you will get it. Or you can draw a tree.

OpenStudy (princeharryyy):

@harz40 = 2*2*2*5, 72 = 2*2*2*3*3 2*2*2 is common n both so, 8 is the answer

OpenStudy (anonymous):

are you sure the answer is 8?

OpenStudy (anonymous):

what do you think ? @UsukiDoll

OpenStudy (usukidoll):

I said either draw a number tree and find out what they have in common or use gcd(40,72) which translates to division. 40 divided by 72. Then you do it again and again until you have gcd(0,1)

OpenStudy (usukidoll):